# Comparing strange and non-strange quark stars within resummed QCD at NLO

### v1

@@ V Numerical Results
Figure 5: Mass-radius relation given by the different approximations. Also included are the NICER data for pulsars PSR J0030+0451 [64, 65] and PSR J0740+6620 [66, 67], as well as the low mass compact star HESS J1731-347 [69]. In particular, the ellipses represent the 68% (dashed) and 95% (full) confidence interval of the 2-D distribution in the mass-radii domain while the error bars give the 1-D marginalized posterior distribution for the same data. A band identifying the mass of the low mass compact object associated with GW190814 [80] has also been included.
@@
With respect to remnant scale dependence, it is important to notice that these values indicate that the RGOPT is moderately more stable to scale variations than pQCD. Indeed, to reproduce the maximum mass variation $\Delta M_{\rm max}=(2.6-2.0)M_{\odot}$ for SQS this method requires a relative scale variation $\Delta X=0.67$ which is slightly higher than the one required by pQCD, $\Delta X=0.61$. A similar situation occurs in the case of NSQS when the OPT requires $\Delta X=0.79$ while pQCD only requires $\Delta X=0.36$. On the different panels of Fig. 4, the dots identify the central chemical potential or baryonic density of the maximum mass configurations. At the same time, the squares on the bottom panel indicate the $P=0$ condition that defines the QS surface. Note that, at the surface, $s$-quarks are already present.
Fig. 5 displays the mass-radius relations obtained with the EoSs of Fig 4 together with the predicted masses and radii of the pulsars PSR J0030+0451 [64, 65] and PSR J0740+6620 [66, 67] as well as the compact object HESS J1731-347 [69]. For reference, a band identifying the low mass compact object associated with GW190814 [80] is also shown. Comparing the results for NSQS with SQS, we find that the NLO RGOPT results show a slight change in the neutron star radii, which are lower for the SQS. This could be expected, since the onset of a new degree of freedom softens the EoS. As a consequence of the softening, in order to describe two solar mass stars or above, larger scale values $X$ are required. In the case of SQM the $X$ values are slightly larger than the ones obtained for pQCD. We note, however, that for the maximum mass range considered [2:2.6]$M_{\odot}$, the NLO RGOPT predictions for both NSQS and SQS properties that are compatible with the present NS observations. More stringent constraints are necessary to distinguish the different scenarios. Considering the choice of scale adopted, the descriptions within both pQCD and RGOPT for $N_{f}=2+1$ are compatible with each other, with NLO RGOPT predicting slightly larger radii. Compared with NNLO pQCD, NLO RGOPT gives a smaller reduction in the radii predicted by SQS (not more than 1 km) compared to NSQS. This is explained by the large difference between the pQCD expressions for $m\neq 0$ and $m=0$, while RGOPT produces similar expressions in both cases, since even in the massless sector, the prescription introduces a resummed medium-dressed mass, $\overline{\eta}(g,\mu)$.
@@ VI Conclusions
[…]
@@ Acknowledgments
T.E.R acknowledges support from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Process SEI-260003/019683/2022. M.B.P. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No 307261/2021-2 and by CAPES - Finance Code 001. T.E.R and M.B.P are also partially supported by Instituto Nacional de Ciência e Tecnologia de Física Nuclear e Aplicações (INCT-FNA), Process No. 464898/2014-5. C.P. received support from Fundação para a Ciência e a Tecnologia (FCT), I.P., Portugal, under the projects UIDB/04564/2020 (doi:10.54499/UIDB/04564/2020), UIDP/04564/2020 (doi:10.54499/UIDP/04564/2020), and 2022.06460.PTDC (doi:10.54499/2022.06460.PTDC).

### v2

@@ V.2 Quark stars
Figure 5: Mass-radius relation given by the different approximations. Lines reproducing maximum QS masses of $2.6M_{\odot}$ correspond to the higher values of $X$ of table 1, while the ones for $2M_{\odot}$ correspond to the lowest $X$ values. Also included are the NICER data for pulsars PSR J0030+0451 [64, 65] and PSR J0740+6620 [66, 67], as well as the low mass compact star HESS J1731-347 [69]. In particular, the ellipses represent the 68% (dashed) and 95% (full) confidence interval of the 2-D distribution in the mass-radii domain while the error bars give the 1-D marginalized posterior distribution for the same data. A band identifying the mass of the low mass compact object associated with GW190814 [82] has also been included.
@@
Having discussed the EOS of $\beta$-equilibrium quark matter, for both SQM and NSQM, we next consider NS observations to constrain the scale $X$. We will apply our EOS to the description of quark stars (QS). We will not consider hybrid stars with a core of quark matter in order to avoid an extra uncertainty through the hadronic EoS.
In Fig. 4 we display the EoSs, pressure versus chemical potential and pressure versus the energy density (top panels) and the respective quark fractions (bottom panel) for the values of $X$ that reproduce $M_{\rm max}=2,\,2.3$ and $2.6M_{\odot}$. The corresponding values  [^9] Notice that the $X$ values of the RGOPT and pQCD for NSQM ($N_{f}=2$) slightly differ from those of Ref. [32] since in that work the scale was chosen to be $\Lambda=X(\mu_{u}+\mu_{d})/2$. for both approximations in the case of SQS and NSQS are given in Table 1. In this table, one can see the NLO RGOPT may require $X$ values slightly higher than those of NNLO pQCD, namely X=3.634.30X=3.63-4.30 for SQS and X=3.394.18X=3.39-4.18 for NSQS, implying that such stars may be composed by weakly interacting quarks ( smaller $\alpha_{s}(X)$), even at low $P$-values. In contrast, the NNLO pQCD produces the required maximum masses for values of $X$ within the canonical range, namely X=2.953.56X=2.95-3.56 for SQS and X=2.172.53X=2.17-2.53 for NSQS. We also checked that using higher orders in the running coupling and mass instead of Eqs.(25),(26) would have very moderate, not much visible impact on our results, which is partly due to our use of exact two-loop running coupling, Eq.(25), less sensitive to higher order differences than the truncated power-expansion in $\ln^{-1}(\Lambda/\mu)$ more commonly used in the literature. Moreover, note that all our RGOPT results in Table 1 lie within a range of $\mu_{B}$ and $X$ values such that the coupling $\alpha_{s}$ remains moderately perturbative, $0.35\leq\alpha_{s}\leq 0.38$ roughly for the values of the scale at the surface of the stars with maximum masses.
We should also comment on the configurations shown in Table 1 by taking into account the stability constraints generally considered for quark matter: i) for stable strange matter, the energy per baryon at the bulk should be below 930 MeV [15, 17], while, for the same input parameters in a given model, $ud$ quark matter should be no more bound than iron, i.e. the energy per baryon should be above 930 MeV, see the discussion in Ref.[83]; ii) for stable $ud$ quark matter, the energy per baryon should be below 930 MeV. Recently, other scenarios have been proposed within phenomenological quark models [19, 29], and therefore the conditions of absolute stability are still under discussion.
Considering first the results obtained with pQCD given in Table 1, they show stable NSQM only for $X\gtrsim 2.5$, corresponding to $M_{\rm{max}}\gtrsim 2.56M_{\odot}$. Looking at the relevant pQCD $X$ values for SQM, the corresponding energies per baryon are below 930 MeV in bulk for almost all values listed in Table 1, i.e. for $X\gtrsim 3$. However, at the same $X$ values, the NSQM energies per baryon are also below 930 MeV, but would give very large corresponding maximum masses $M_{\rm{max}}\gtrsim 2.7M_{\odot}$. In several recent studies [84, 85] it is shown, using different agnostic descriptions of the neutron star EOS and imposing pQCD constraints, that the maximum mass of the NS is below $\sim 2.5M\odot$. Therefore, according to these studies models with $M_{\rm{max}}\geq 2.7M_{\odot}$ appear unfavored. However, some works propose the existence of quark stars with masses above $2.5M\odot$: in Ref. [86] a scenario of uds QS with a maximum mass of $2.5-2.6M_{\odot}$ is associated to the binary neutron star merger GW190814, and in Ref. [28] the authors obtain within an interacting quark model $M_{\rm{max}}\leq 3.23M_{\odot}$, a value of the order proposed by Rhoades and Ruffini [87] as maximum NS mass determined just from the Einstein’s theory of relativity, the principle of causality, and Le Chatelier’s principle. Moreover, note that pQCD constraints are treated differently in Ref. [70], where quark stars are discussed.
@@
The two scenarios obtained with both approaches (pQCD or RGOPT) for stable NSQM do not appear to be realistic according to several different studies that propose maximum masses of the order of 2.15-2.35 $M_{\odot}$ [88, 89, 90, 84] and generally below 2.5 $M_{\odot}$ [84]. Therefore, within the possible limitations of our level of approximations, and considering realistic maximum masses between 2.3 and 2.5 solar masses, our descriptions seem to agree with the Bodmer-Witten conjecture on SQM, with $ud$ quark matter being less bound than iron. This could suggest that pure quark stars would be made of udsuds matter, while $ud$ quark stars would carry an outer layer of hadronic matter or could be quark stars. These results indicate that the composition of dense matter is still not clear and further studies are required.
With respect to remnant scale dependence, it is important to notice that these values indicate that the RGOPT is moderately more stable to scale variations than pQCD. Indeed, to reproduce the maximum mass variation $\Delta M_{\rm max}=(2.6-2.0)M_{\odot}$ for SQS this method requires a relative scale variation $\Delta X=0.67$, which is slightly higher than the one required by pQCD, $\Delta X=0.61$. A similar situation occurs in the case of NSQS when the OPT requires $\Delta X=0.79$ while pQCD only requires $\Delta X=0.36$. On the different panels of Fig. 4, the dots identify the central chemical potential or baryonic density of the maximum mass configurations. At the same time, the squares on the bottom panel indicate the $P=0$ condition that defines the QS surface. Note that on the surface $s$-quarks are already present.
[…]
@@ VI Conclusions
After carrying out those modifications, in accordance with Ref. [32], we have determined the EoS representing the two scenarios, with and without the $s$-quark. In more academic applications, the value of the parameter characterizing the scale, is generally chosen conventionally to be $X\in[1,4]$ with $X=2$ representing the so-called central scale. In the present work, this arbitrary scale parameter was fixed in a more phenomenological way by selecting the values which give maximum QS masses of order $M_{\rm max}=2,\,2.3$ and $2.6\,M_{\odot}$ which represent some of the limits often considered in the literature. For the NLO RGOPT EoS the required values are, respectively, X=3.63, 3.97X=3.63,\,3.97 and  4.30\,4.30 when $N_{f}=2+1$ and X=3.39, 3.80X=3.39,\,3.80 and $4.18$ when $N_{f}=2$ whereas for the NNLO pQCD EoS the values are X=2.95, 3.26X=2.95,\,3.26 and $3.56$ when $N_{f}=2+1$ and X=2.17, 2.35X=2.17,\,2.35 and $2.53$ when $N_{f}=2$. Accordingly, within the RGOPT framework, massive QS can only be produced at relatively larger scales (smaller couplings) compared with pQCD. The introduction of an extra degree of freedom, the $s$-quark, softens the EoS and to attain the same maximum masses the $N_{f}=2+1$ scenario requires larger $X$ scales in both descriptions. With these choice of the scale, it was shown that the EoS obtained are in accordance with properties of several recent neutron star mass-radius data, in particular, with data for pulsars PSR J0030+0451 [64, 65] and PSR J0740+6620 [66, 67], and for the low mass compact star HESS J1731-347. To understand these results, let us first recall that the introduction of finite quark masses tends to reduce the pressure predicted by both methods, as discussed earlier. This effect is amplified in the RGOPT approach, since the effective quark masses are $m_{s}+\overline{\eta}$ (in the strange heavy sector) and $\overline{\eta}$ (in the non-strange light sector), implying that the corresponding EoS for the same scale $X$ and baryonic chemical potential reach lower pressures than their pQCD counterparts. Therefore, the RGOPT predicts that QS with high masses can only be formed when the scale is rather high. Comparing the $N_{f}=2+1$ and $N_{f}=2$ RGOPT mass-radii curves one can see that, despite their similarities and good agreement with observations, the description of strange QS requires $X>3.63$ implying that the coupling takes relatively lower numerical values. The $_{f}=2+1 NN$LO pQCD description of SQS also calls for somewhat higher $X$ values which, however, lie within the more canonical range. Note also that the RGOPT is slightly more stable to scale variations than pQCD. In particular, to reproduce the maximum mass variation $\Delta M_{\rm max}=(2.6-2.0)M_{\odot}$ corresponding to possible different scenarios for SQS, NLO RGOPT requires a relative scale variation of $\Delta X=0.67$, about $10\%$ larger than the variation occurring in pQCD, $\Delta X=0.61$. We have also studied more precisely the stability of SQM and NSQM scenarios upon examining the energy per baryon for our different QM approximations. For both pQCD and RGOPT, $ud$ quark matter appear less bound than iron, except for very large masses $M_{\rm{max}}\gtrsim 2.6M_{\odot}$. On the other hand, the NLO RGOPT predictions are in agreement with the SQM Bodmer-Witten conjecture within the mass range $2.30M_{\odot}\lesssim M_{\rm{max}}\lesssim 2.54M_{\odot}$. At the same time, they satisfy the condition that $ud$ matter is less bound than iron. Note, that it is difficult to assert firm conclusions about Bodmer-Witten conjecture without a truly non-perturbative contribution to the QCD vacuum energy, that neither pQCD nor RGOPT incorporate.
@@ Acknowledgments
T.E.R acknowledges support from Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Process SEI-260003/019683/2022. M.B.P. is partially supported by Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Process No. 307261/2021-2 and 403016/2024-0. T.E.R and M.B.P are also partially supported by Instituto Nacional de Ciência e Tecnologia de Física Nuclear e Aplicações (INCT-FNA), Process No. 464898/2014-5. C.P. received support from Fundação para a Ciência e a Tecnologia (FCT), I.P., Portugal, under the projects UIDB/04564/2020 (doi:10.54499/UIDB/04564/2020), UIDP/04564/2020 (doi:10.54499/UIDP/04564/2020), and 2022.06460.PTDC (doi:10.54499/2022.06460.PTDC).